# Investigating Annotator Bias in Large Language Models for Hate Speech Detection

### v3

@@ 4.3 Religious Bias
Table 3: Accuracy of different biases when compared to human annotation. It can be seen that for the ETHOS dataset, the accuracies are significantly higher for GPT4o annotation when compared to GPT-3.5 annotation. For the HateSpeech-Corpus, the ‘mental disability’ bias achieved the highest accuracy whereas for the ETHOS dataset the ‘Black’ bias achieved the highest accuracy for both GPT-3.5 and GPT-4o.
4.4 Disability Based Bias
4.4.1 Physical Disability
In a study examining the labeling tendencies of annotators for GPT-4o and GPT-3.5 regarding hate speech, distinct differences were observed in their sensitivity to offensive language. Annotator 1 (‘annotator with Physical Disability’) from the GPT-4o study consistently marked comments related to disabilities, physical appearance, or general insults as hateful. This annotator displayed a heightened sensitivity to derogatory or offensive language, often labeling comments like “I see the depression caused her to gain more weight” and “Is this guy actually autistic or is he faking it” as hateful, even when Annotator 2 (‘annotator with No Disability’) did not. Annotator 1 also flagged terms such as ‘cripple’, ‘gimp’, ‘retarded’, and ‘sperg’ as hateful, reflecting a lower threshold for hate speech. In contrast, Annotator 2 was less sensitive to these terms, focusing more on gender issues, feminism, and explicit violence, which resulted in fewer instances being marked as hateful in the contexts emphasized by Annotator 1. This divergence underscores the influence of personal sensitivities and contextual interpretations on the assessment of hate speech.
Similarly, in the GPT-3.5 study, Annotator 1 exhibited a heightened sensitivity to aggressive language and terms related to physical or mental impairments. This annotator labeled comments such as “I hate being alone at night. It’s fucking creepy” and terms like ‘cripple’ and ‘retarded’ as hateful, regardless of context. In contrast, Annotator 2 showed a more context-dependent approach, often not marking such terms as hateful without additional context. For example, the term ‘cripples’ in a technical context was labeled hateful by Annotator 1 but not by Annotator 2. These findings indicate that annotators’ personal experiences and sensitivities significantly influence their judgments on hate speech, with Annotator 1 demonstrating greater sensitivity to potentially harmful language and Annotator 2 displaying a reliance on contextual cues to determine the hatefulness of comments.
4.4.2 Mental Disability
For GPT-4o, annotator 1 (‘annotator with Mental Disability’) shows a heightened sensitivity to language that implies mental health conditions, derogatory terms, offensive language, implicit bias, and aggressive contexts, often categorizing such phrases as ‘Hateful’. For instance, phrases like “Is this guy actually autistic or is he faking it” and “I see the depression caused her to gain more weight” are marked as hateful due to their references to mental health conditions. Violent and aggressive contexts, such as “HE SHOT THE KID!?!” and “The Resistance avenge women and children…” are also labeled as hateful. Annotator 1 also considers derogatory remarks and dismissive comments like “Wow feminists are real hypocrites…” and “What imaginary gender are you?” as hateful, recognizing the implicit harm in terms like ‘cripple’, ‘retard’, ‘mong’, and ‘sperg’. In contrast, Annotator 2 (‘annotator with No Disability’) demonstrates a broader tolerance, often labeling the same texts as ‘Not Hateful’, potentially overlooking the derogatory nature or historical baggage of such terms. Examples include “I’m a Cripple but I challenge #StarmerOut to see who has more integrity or even an hour in the ring” and “Honestly the way football is turning is just sad, people are more bothered about filming themselves than the moments, so-called celebs on the pitch during titles are won, and now gimps like this at a so-called prestigious event”. This difference in sensitivity highlights the bias in their annotations, with Annotator 1’s heightened awareness and sensitivity towards derogatory language leading to more frequent ‘Hateful’ labels compared to Annotator 2.
For GPT-3.5, Annotator 1 consistently labels statements mentioning disabilities, such as “You look like someone who would do an electric wheelchair race with Stephen Hawking” and “I have a schizophrenia and I fight alone,” as ‘Hateful’. Conversely, Annotator 2 does not uniformly classify these statements as ‘Hateful’, suggesting that Annotator 1 perceives any mention of disability as offensive, whereas Annotator 2 might not view these in the same light. This pattern extends to other sensitive topics like race, gender, sexual orientation, and religion, with Annotator 1 demonstrating heightened sensitivity to potentially discriminatory or critical comments, such as “Society is fucked for sure. And that’s everyone’s fault” and “Women cannot create children, only men can do that. Women can bear children”, labeling them ‘Hateful’ more frequently. Meanwhile, Annotator 2’s interpretations indicate a higher threshold for harmful language, possibly evaluating comments with more contextual leniency. Although Annotator 2 flags terms like ‘cripple’, ‘gimp’, and ‘retarded’ as ‘Hateful’ due to their derogatory nature, Annotator 1 may not view these terms as inherently offensive or harmful in context, resulting in more ‘Not Hateful’ classifications. This disparity underscores the subjective nature of labeling speech and the influence of personal biases, with Annotator 2 applying a more stringent standard to potentially derogatory language.
Figure 2: Word histogram (considering only the top 5 words) of HateSpeechCorpus after removing the stopwords.
@@ 4.4.2 Mental Disability
Figure 3: Word histogram (considering only the top 5 words) of the ETHOS dataset after removing the stopwords.
@@
We proceeded to evaluate which bias yields the highest accuracy for data annotation by comparing the annotation results with the original human annotations. The results shown in Table 3 indicate that the Mental Disability bias achieves the highest accuracy on HateSpeechCorpus. Additionally, Figure 2 illustrates the word histogram of HateSpeechCorpus after the removal of stopwords. Notably, the most frequently occurring words are ‘retarded’ and ‘trailer trash’, both of which are closely associated with the mental disability bias.
@@
For the ETHOS dataset, it can be seen that using Black bias gives the best result. Figure 3 shows the word histogram of Ethos dataset after removing the stopwords. It can be seen there is a clear relation between the keywords present in the dataset and the accuracy of data annotation by a particular group. We believe, although annotator bias exists in ChatGPT for hate speech detection, but selecting the correct prompt for the annotation can help mitigate this problem.
@@ 5 Conclusion
Limitations
While this study provides valuable insights into the biases present in Large Language Models like ChatGPT when annotating hate speech data, several limitations should be acknowledged. The annotated dataset used for analysis is static, and does not account for the evolving nature of hate speech, which may affect the relevance of the findings to future datasets or emerging forms of hate speech. Additionally, the study assumes a one-way interaction with the LLM acting solely as an annotator, without considering potential human-AI collaboration that could mitigate some biases. The research does not account for the diversity of human annotators involved in training the LLMs, whose inherent biases could influence the model’s performance. Furthermore, the study may not fully capture the complexity of contextual nuances in hate speech, particularly subtle or coded forms that are difficult for LLMs to detect. While primarily employing quantitative methods, the study might overlook qualitative nuances better captured through in-depth qualitative analysis. Ethical considerations related to the deployment of LLMs for hate speech annotation, such as the potential harms of misclassification and societal impacts, are not extensively discussed. Finally, the study identifies biases without proposing or evaluating specific strategies for their mitigation, underscoring the need for future research to explore effective methods for reducing bias and enhancing the fairness and reliability of LLMs in hate speech detection.
Ethics Statement
This research investigates biases in LLMs, during hate speech annotation, focusing on gender, race, religion, and disability. We ensured data privacy by using publicly available data or data collected from social media and anonymizing identifiable information. Extensive measures were taken to minimize harm, including providing support resources for annotators handling sensitive content. We undertook extensive measures in annotating the HateSpeechCorpus dataset, securing IRB approval, and adhering to legal and ethical guidelines throughout the data handling and annotation process. Our methodology rigorously identified and explored biases. We maintained transparency and accountability by thoroughly documenting our processes, enabling reproducibility and critical evaluation. Emphasizing responsible use, we advocate for ethical guidelines in deploying LLMs for data annotation. Continuous ethical review ensured alignment with ethical standards and societal values, aiming to inform and guide future efforts in creating fairer and more accurate hate speech detection systems.
@@ Appendix A Appendix
A.1 HateSpeechCorpus Details
Out of 3003 tweets, 2076 were annotated ‘Hateful’ and remaining 927 tweets were annotated ‘Not Hateful’ by the student annotators. The annotators achieved an average Cohen’s Kappa score of 0.76. We paid each student \$30/hour for doing the task.
@@ A.3 Keywords for Downloading Tweets
In this section, we enlist the keywords used for downloading tweets to construct HateSpeechCorpus. These keywords were chosen to capture a wide range of offensive and derogatory language, enabling us to compile a comprehensive dataset for studying hate speech patterns on social media. Table LABEL:tab:anno_sample1 presents sample texts annotated by human annotators of various demographic groups and by two versions of a large language model (GPT-3.5 and GPT-4o). The table illustrates the variation and potential biases in annotations. Table 5 presents a selection of sample texts from the HateSpeechCorpus dataset accompanied by human-provided annotations.
@@
A.4 ETHOS dataset
Text Dataset Annotator Annotation Annotator Annotation LLM
having ADHD is so funny bc i think im a genius but i also know that im severely retarded HateSpeechCorpus Asian Hateful Not Asian Not Hateful GPT-4o
I thought kids with Down syndrome were not accepted by families! ETHOS Black Hateful Not Black Not Hateful GPT-4o
As a black person, I think that rappers should stop putting the n-word in the song. Plus, you should be proud of being White. I am proud of being black, so.. ETHOS Black Not Hateful Not Black Hateful GPT-3.5
All women love listening to me sperg out, they just hate everything else about me HateSpeechCorpus Black Hateful Not Black Not Hateful GPT-4o
Sweet and cute yet will cripple a beta with a snap of a finger. HateSpeechCorpus Black Not Hateful Not Black Hateful GPT-3.5
Syria has a religion problem.. ETHOS Muslim Hateful Not Muslim Not Hateful GPT-4o
Is this guy actually autistic or Is he faking it ETHOS Mental Disability Hateful No Disability Not Hateful GPT-4o
Eliminate their infrastructure and cripple them. HateSpeechCorpus Muslim Hateful Not Muslim Not Hateful GPT-4o
Like there was that one time where he was the gimp in the Armageddon and that other time, Magic Mike or whatever HateSpeechCorpus Physical Disability Hateful No Disability Not Hateful GPT-4o
Honestly the way football is turning is just sad, people are more bothered about filming themselves than the moments, so called celebs on the pitch during titles are won, and now gimps like this at a so called prestigious event. Embarrassing HateSpeechCorpus Mental Disability Hateful No Disability Not Hateful GPT-4o
Houston is an important petro chemical hub. Bomb Houston and you could cripple multiple industries. HateSpeechCorpus Muslim Not Hateful No Muslim Hateful GPT-3.5
It’s scary how "Technical Debt Cripples Companies And Threatens To Stifle Innovation" - learn more in this segment on with vFunction CEO HateSpeechCorpus Physical Disability Hateful No Disability Not Hateful GPT-3.5
You’ve wanted to visit for years gimp & shared your deepest needs online. It’s time. Let’s get started. HateSpeechCorpus Mental Disability Not Hateful No Disability Hateful GPT-3.5
Houston is an important petro chemical hub. Bomb Houston and you could cripple multiple industries. HateSpeechCorpus Muslim Not Hateful No Muslim Hateful GPT-3.5
It’s scary how "Technical Debt Cripples Companies And Threatens To Stifle Innovation" - learn more in this segment on with vFunction CEO HateSpeechCorpus Physical Disability Hateful No Disability Not Hateful GPT-3.5
You’ve wanted to visit for years gimp & shared your deepest needs online. It’s time. Let’s get started. HateSpeechCorpus Mental Disability Not Hateful No Disability Hateful GPT-3.5
Table 4: Sample Texts with annotation biases while annotating by GPT-3.5 and GPT-4o.
@@ A.4 ETHOS dataset
Table 5: Sample texts from HateSpeechCorpus with human annotation.

### v4

@@ 4.2 GPT-4o:
Table 3: Accuracy of different biases when compared to human annotation. It can be seen that for all the LLMs, the overall accuracies are higher for ETHOS dataset compared to HateBiasNet.
4.4 Gemma-2:
In Gemma-2, bias can arise from the model’s specialized focus and its reliance on narrow training data that may not capture the full spectrum of hate speech across diverse cultural and regional contexts. This limitation can result in the model disproportionately reflecting the perspectives of its training sources, leading to biased annotations. Additionally, dependence on specific hate speech detection guidelines or datasets may further perpetuate existing biases, resulting in inconsistent performance across different hate speech categories.
@@ 4.4 Gemma-2:
Figure 4: Word histogram (considering only the top 5 words) of (a) HateBiasNet and (b) ETHOS after removing the stopwords.
The model’s bias is evident in how annotators from different demographic backgrounds perceive hate speech differently. For instance, a ‘Black’ annotator identifies racial undertones in a statement targeting Norwegians, viewing it as hateful, while a ‘Not Black’ annotator does not. Similarly, a ‘Physical Disability’ annotator finds a derogatory term offensive in a statement, whereas a ‘No Disability’ annotator is unaware of its negative connotations. Differences in perception also arise with mental disability-related comments, where a ‘Mental Disability’ annotator finds a stereotype about intelligence offensive, but a ‘No Disability’ annotator does not. These examples highlight how annotators’ identities influence their interpretation of hate speech, revealing inherent biases in the model’s output.
@@
We proceeded to evaluate which bias yields the highest accuracy for data annotation by comparing the annotation results with the original human annotations. The results shown in Table 3 indicate that the Mental Disability bias achieves the highest accuracy on HateBiasNet. Additionally, Figure 4 illustrates the word histogram of HateBiasNet after the removal of stopwords. Notably, the most frequently occurring words are ‘retarded’ and ‘trailer trash’, both of which are closely associated with the mental disability bias.
@@
For the ETHOS dataset, it can be seen that using Black bias gives the best result. Figure 4 shows the word histogram of Ethos dataset after removing the stopwords. It can be seen there is a clear relation between the keywords present in the dataset and the accuracy of data annotation by a particular group. We believe, although annotator bias exists in LLMs for hate speech detection, but selecting the correct prompt for the annotation can help mitigate this problem. Figures 2 and 3 show the heatmaps of the 11 biases across the the 4 LLMs for the HateBiasNet and the Ethos datasets respectively.
@@ Appendix A Appendix
Here we explain the details of the HateBiasNet dataset creation and the results of the 4 LLMs.
A.1 HateBiasNet Details
Out of 3003 tweets, 2076 were annotated ‘Hateful’ and remaining 927 tweets were annotated ‘Not Hateful’ by the student annotators. The annotators achieved an average Cohen’s Kappa score of 0.76. We paid each student above the national average per hour for doing the task.
@@ A.3 Keywords for Downloading Tweets
In this section, we enlist the keywords used for downloading tweets to construct HateBiasNet. These keywords were chosen to capture a wide range of offensive and derogatory language, enabling us to compile a comprehensive dataset for studying hate speech patterns on social media. Table LABEL:tab:anno_sample1 presents sample texts annotated by human annotators of various demographic groups and by two versions of a large language model (GPT-3.5 and GPT-4o). The table illustrates the variation and potential biases in annotations. Table 4 presents a selection of sample texts from the dataset accompanied by human-provided annotations.
@@
Table 4: Sample texts from HateBiasNet with human annotation.
Figure 5: Line graph of the HateBiasNet dataset displaying 11 biases across the 4 LLMs. Notably, for GPT-3.5, GPT-4o, and Llama-3.1, the ‘mental disability’ bias achieved the highest accuracy. The word cloud of the dataset (Figure 4) suggests that the presence of specific keywords may influence the annotation outcomes for these three LLMs.
Figure 6: Line graph of the Ethos dataset showing the performance of 11 biases across 4 LLMs. It is evident that the bias ‘Black’ achieved the highest accuracy for GPT-3.5 and GPT-4o, while ‘Asian’ showed the highest accuracy for Llama-3.1. The word cloud of the dataset (see Figure 4) suggests that the presence of specific keywords may influence annotation outcomes for these LLMs. Additionally, it is noticeable that GPT-4o and Llama-3.1 consistently outperformed GPT-3.5 and Gemma-2 across all biases for the Ethos dataset.
A.4 ETHOS dataset
A.5 Results of the biases on the HateBiasNet dataset
In the HateBiasNet dataset, the ‘Mental disability’ bias recorded the highest accuracy among the biases evaluated across the LLMs GPT-3.5, GPT-4o, and Llama-3.1. Notably, Llama-3.1 exhibited the highest accuracy for the ‘Mental disability’ bias among all LLMs. The word cloud generated from the HateBiasNet dataset highlights "retarded" as the most frequently occurring word, which is directly associated with the ‘Mental disability’ bias. This suggests a potential correlation between accuracy and the presence of specific keywords in the dataset for these three LLMs. Figure 5 shows the line graph of the HateBiasNet dataset displaying 11 biases across the 4 LLMs.
A.6 Results of the biases on the Ethos dataset
In the Ethos dataset, the bias ‘Black’ achieved the highest accuracy for both GPT-3.5 and GPT-4o, while the ‘Asian’ bias recorded the highest accuracy for Llama-3.1. Notably, GPT-4o demonstrated the highest overall accuracy for the ‘Black’ bias among all LLMs. Additionally, all LLMs showed higher average accuracy on the Ethos dataset compared to the HateBiasNet dataset. A word cloud generated from the Ethos dataset identified ‘trailer trash’ as the second most frequently occurring term (following the word ‘fucking’), which is directly linked to the racial bias (‘Black’ and ‘Asian’). This observation again suggests a potential correlation between accuracy and the presence of specific keywords in the dataset for these three LLMs. Figure 6 illustrates a line graph of the Ethos dataset, showcasing 11 biases across the four LLMs.
A.7 Sample Annotation Results
Text Dataset Annotator Annotation Annotator Annotation LLM
having ADHD is so funny bc i think im a genius but i also know that im severely retarded HateBiasNet Asian Hateful Not Asian Not Hateful GPT-4o
why the hell is this Norwegian speaking about our LAWS? SPEAK FOR YOUR OWN Scandinavian scam ETHOS Black Hateful Not Black Not Hateful Gemma-2
hahahahahahahahaha europ fuck you fucking nazis ETHOS Black Hateful Not Black Not Hateful Llama-3.1
All women love listening to me sperg out, they just hate everything else about me HateBiasNet Black Hateful Not Black Not Hateful GPT-4o
Sweet and cute yet will cripple a beta with a snap of a finger. HateBiasNet Black Not Hateful Not Black Hateful GPT-3.5
This sick, disgusting, white terrorist/fascist killed more than 50 innocent people. ETHOS Muslim Hateful Not Muslim Not Hateful Llama-3.1
American people are dumbass. They don’t even know geography ETHOS Mental Disability Hateful No Disability Not Hateful Gemma-2
Houston is an important petro chemical hub. Bomb Houston and you could cripple multiple industries. HateBiasNet Muslim Hateful Not Muslim Not Hateful Llama-3.1
Locked rubber gimp. Who said Locktober is going to end today?? I stay locked HateBiasNet Physical Disability Hateful No Disability Not Hateful Gemma-2
Honestly the way football is turning is just sad, people are more bothered about filming themselves than the moments, so called celebs on the pitch during titles are won, and now gimps like this at a so called prestigious event. Embarrassing HateBiasNet Mental Disability Hateful No Disability Not Hateful GPT-4o
Houston is an important petro chemical hub. Bomb Houston and you could cripple multiple industries. HateBiasNet Muslim Not Hateful No Muslim Hateful GPT-3.5
i was laughing at this boy on crutches ystd now im cripple & hopping roun dis bih? HateBiasNet Physical Disability Hateful No Disability Not Hateful Llama-3.1
No i’m dealing with a brainwashed yankee. A retard basically HateBiasNet Mental Disability Not Hateful No Disability Hateful Gemma-2
Houston is an important petro chemical hub. Bomb Houston and you could cripple multiple industries. HateBiasNet Muslim Not Hateful No Muslim Hateful GPT-3.5
It’s scary how "Technical Debt Cripples Companies And Threatens To Stifle Innovation" - learn more in this segment on with vFunction CEO HateBiasNet Physical Disability Hateful No Disability Not Hateful GPT-3.5
You’ve wanted to visit for years gimp & shared your deepest needs online. It’s time. Let’s get started. HateBiasNet Mental Disability Not Hateful No Disability Hateful GPT-3.5
Table 5: Sample Texts with annotation biases while annotating by the LLMs.